# Emergence of tunable exceptional points in altermagnet-ferromagnet junctions

### v1

@@ Abstract
Introduction– Non-Hermitian (NH) systems have gained extensive attention in both theoretical and experimental research in recent years Moiseyev (2011); Bender (2007); Ashida et al. (2020); Bergholtz et al. (2021); Kawabata et al. (2019); Banerjee et al. (2023a). This growing interest stems from the realization of intriguing NH phenomena, such as EPs, the NH skin effect, exotic topological phases, generalized Bloch band theory, and extended symmetry classes, to name just a few. EPs, where both eigenvalues and eigenstates coalesce, are fundamentally unique to NH physics Heiss (2012); Kato (2013). EPs are associated with a plethora of fascinating phenomena, both in theory and applications, including NH topological phases linked with winding of eigenvalues and eigenvectors Ding et al. (2022); Banerjee et al. (2023b), enhanced sensing Hodaei et al. (2017), optical microcavities Chen et al. (2017), and directional lasing technologies Peng et al. (2016). EPs have been realized across diverse experimental platforms, encompassing fields such as optics, photonics, electric circuits, and acoustics Parto et al. (2020); Miri and Alù (2019); Stehmann et al. (2004); Choi et al. (2018); Zhu et al. (2018); Shi et al. (2016); Zhu et al. (2018). In addition, they have been theoretically investigated at material interfaces, including topological insulator-FM Bergholtz and Budich (2019) and superconductor-FMCayao and Black-Schaffer (2022) junctions.
@@
On the other hand, AMs are a newly-identified magnetic class distinct from conventional FMs and antiferromagnets Šmejkal et al. (2022a, b). In AMs, opposite spin sublattices are linked by a rotational symmetry rather than the conventional translational or inversion operations Šmejkal et al. (2022b); Litvin and Opechowski (1974); Brinkman and Elliott (1966), which results in a subtle interplay of characteristics of both FMs and antiferromagnets. They demonstrate spin-split band structures, similar to FMs, but have zero total magnetization akin to antiferromagnets Šmejkal et al. (2022b, 2020). The library of altermagnetic materials has been expanding rapidly Šmejkal et al. (2022b); Fedchenko et al. (2024); Šmejkal et al. (2022a); Yuan et al. (2020); Šmejkal et al. (2022a, b); Krempaskỳ et al. (2024); Zhu et al. (2024); Devaraj et al. (2024). Furthermore, AM junctions have been recently investigated for spin pumping Sun and Linder (2023) and intriguing transport properties Das et al. (2023); Lyu and Li (2024); Papaj (2023); Ouassou et al. (2023); Beenakker and Vakhtel (2023); Niu and Zhang (2024); Giil and Linder (2024); Cheng et al. (2024); Das and Soori (2024); Sukhachov et al. (2024); Niu and Yang (2024); Lu et al. (2024); Banerjee and Scheurer (2024); Nagae et al. (2024).
@@
Setup of AM-FM junctions– We consider an AM-FM junction by coupling a two-dimensional AM with a semi-infinite FM lead as depicted in Fig. 1. Region $z<0$ is the FM lead and $z=0$ is the interface. We treat the AM-FM junction as an open quantum system and model it with an effective Hamiltonian of the form,
@@
where $H_{\text{AM}}$ denotes the Hamiltonian of the considered AM system (discussed later), and $\Sigma_{L}$ represents the self-energy arising from the semi-infinite FM lead. In the wide-band limit, the self-energy term becomes independent of both momentum and frequency, and can be analytically calculated as Ryndyk et al. (2009); Datta (1997); Bergholtz and Budich (2019); Cayao and Black-Schaffer (2022),
@@
EPs in d-wave AM– Let us consider the setup with an FM lead attached to a $d$-wave AM, whose Fermi surface is oriented as shown in Fig. 1(a). The Hamiltonian for a $d$-wave AM is given by Šmejkal et al. (2022b)
@@
$\gamma^{2}=\lambda^{2}(k_{x}^{2}+k_{y}^{2})+4t_{j1}k_{x}^{2}k_{y}^{2},\quad\gamma k_{x}k_{y}=0.$ (5)
@@
For the remainder of the paper, for simplicity, we will set $t=1$, $t_{j1}=1$, $\gamma=1$, and $\lambda=1$. The real and imaginary parts of the eigenvalues are plotted in Fig. 2 (a) and (b), respectively. The red dots indicate the merging of the eigenvalues at $(0,\pm 1),(\pm 1,0)$. We further confirm the coalescing of the eigenvectors at the same points by calculating the phase rigidity, $r=\frac{\langle\Psi_{L}|\Psi_{R}\rangle}{\langle\Psi_{R}|\Psi_{R}\rangle}$ Heiss (2012). Here $\Psi_{L}$ and $\Psi_{R}$ denote the left and right eigenvectors respectively. Due to bi-orthogonalization, $r\rightarrow 0$ near EPs and approaches unity away from them. In Fig. 2(c) and (d), we show $r$ along $k_{x}$ and $k_{y}$, respectively. We note that $r$ vanishes at $(0,\pm 1)$ and (±$(\pm 1,0)$, which confirms that not only the eigenvalues merge but also the eigenstates coalesce.
As another characterization of the emergence of EPs, we calculate the vorticity, $\nu_{mn}$, given by Shen et al. (2018)
@@
Here $E_{m}(\mathbf{k})$ and $E_{n}(\mathbf{k})$ denote the energy eigenvalues in the complex-energy plane and $C$ is a closed loop in the momentum space. In Fig. 2(e), we plot the vorticity of one of the EPs [located at $(0,+1)$] by choosing $C$ around it. We clearly observe the braiding of bands, a direct signature of the underlying EP. This further confirms the appearance of EPs. To examine the order of EPs, we plot the phase rigidity in Fig. 2(f). The blue dots indicate the calculated values of $r$, while the solid green line is the fitted function. The fitting revealed that the exponent of scaling of $r$ is $\sim$0.47, which is close to that expected for a second-order EP Bulgakov et al. (2006); Rotter (2009). Having seen the emergence of EPs in AM-FM junctions, next, we will show how applying a magnetic field allows us to tune and control them.
@@
Magnetic field tuning of EPs– We next propose and show that the positions of the EPs can be controlled by applying an external magnetic field, $B$, along the interface plane. We find that $B$ can not only change the EP positions asymmetrically, but also merge them beyond a critical value. For illustration, we include the external magnetic field, in our $d$-wave AM-FM junction, along $x$ direction, as shown in Fig. 1. Then, the Hamiltonian in Eq. 3 is augmented by a term of the form $B\sigma_{x}$. The conditions for EPs now read,
$\gamma^{2}=\lambda^{2}k_{x}^{2}+(B+\lambda k_{y})^{2}+4t_{j1}k_{x}^{2}k_{y}^{2},\quad\gamma k_{x}k_{y}=0.$ (7)
@@
We observe that the position of the EPs can not only be tuned along $k_{x}$ and $k_{y}$ (using $B$), but also can be rotated about the origin by choosing different orientations of the AM Fermi surface with respect to the FM lead. For instance, a $\pi/4$ rotation is achieved with the AM Fermi surface shown in Fig. 1(b). This can be obtained by choosing different AM-specific spin-dependent hopping terms in the Hamiltonian (for details see supplement sup ). We, furthermore, discovered that the position of the EPs can be rotated by an arbitrary angle $\theta$, determined by $\tan\theta=\frac{-t_{j1}\pm\sqrt{t_{j1}^{2}+4t_{j2}^{2}}}{-2t_{j2}}$. Here $t_{j1}$ and $t_{j2}$ are the AM-specific hopping parameters related to the orientations of the Fermi surface. We present detailed calculations in the supplement sup .
@@
EPs in $g$- and $i$-wave AMs– We further extend our analysis to AMs with higher harmonic spin-split Fermi surfaces. Let us first consider a $g$-wave AM-FM junction. The Hamiltonian for a $g$-wave AM reads Šmejkal et al. (2022b)
@@
Strikingly distinct from the $d$-wave case, we find that a total of eight EPs emerge at such a junction. The location of EPs are given by $(0,\pm\frac{\gamma}{\lambda})$, (±$(\pm\frac{\gamma}{\lambda},0)$, (±$(\pm\frac{\gamma}{\sqrt{2}\lambda},\pm\frac{\gamma}{\sqrt{2}\lambda})$. These are shown by red dots in Fig. 5(a) and (b). We can extend this analysis to $i$-wave AMs (see supplement material for details of the Hamiltonian and location of EPs sup ). In this case we find twelve EPs emerging at the AM-FM junction. The real and imaginary parts of the energy eigenvalues are presented in Fig. 5(c) and (d), respectively. We note that, in general, there appear twice the EPs as the spin group integer of the AM Šmejkal et al. (2022a). We further note that an applied magnetic field should allow us to tune the EPs in these $g$- and $i$-wave AM-FM junctions, analogous to the $d$-wave case.
@@
Summary and outlook– We have proposed and demonstrated the emergence of NH physics, particularly EPs, in AM-FM junctions. We have revealed the presence of EPs in $d$-wave AM-FM junctions and shown their tunability using an external magnetic field. Our findings reveal that EPs can annihilate at a critical value of $B$. Additionally, we have found the tunability of EPs by varying the orientation of the AM Fermi surfaces. We discovered that the number of emergent EPs depends on the spin group integer of the underlying AM. Here we have focused on the EPs in AM-based NH systems. In future, it could be worth exploring other NH aspects of altermagnetic systems, including NH skin effects, NH topology, and interplay with various types of dissipation mechanisms.
@@
Acknowledgments– We thank A. Banerjee, A. Bose, and S. Debadatta for useful discussions. M.A.R. is supported by a graduate fellowship of the Indian Institute of Science. A.N. acknowledges support from the DST MATRICS grant (MTR/2023/000021).

### v2

@@ Abstract
I Introduction
Non-Hermitian (NH) systems have gained extensive attention in both theoretical and experimental research in recent years [1, 2, 3, 4, 5, 6]. This growing interest stems from the realization of intriguing NH phenomena, such as EPs, the NH skin effect, exotic topological phases, generalized Bloch band theory, and extended symmetry classes, to name just a few. EPs, where both eigenvalues and eigenstates coalesce, are fundamentally unique to NH physics [7, 8]. EPs are associated with a plethora of fascinating phenomena, both in theory and applications, including NH topological phases linked with winding of eigenvalues and eigenvectors [9, 10, 11], enhanced sensing [12], optical microcavities [13], and directional lasing technologies [14]. EPs have been realized across diverse experimental platforms, encompassing fields such as optics, photonics, electric circuits, and acoustics [15, 16, 17, 18, 19, 20, 19]. In addition, they have been theoretically investigated at material interfaces, including topological insulator-FM [21] and superconductor-FM[22] junctions.
@@ I Introduction
On the other hand, AMs are a newly-identified magnetic class distinct from conventional FMs and antiferromagnets [23, 24]. In AMs, opposite spin sublattices are linked by a rotational symmetry rather than the conventional translational or inversion operations [24, 25, 26], which results in a subtle interplay of characteristics of both FMs and antiferromagnets. They demonstrate spin-split band structures, similar to FMs, but have zero total magnetization akin to antiferromagnets [24, 27]. The library of altermagnetic materials has been expanding rapidly [24, 28, 23, 29, 23, 24, 30, 31, 32]. Furthermore, AM junctions have been recently investigated for spin pumping [33] and intriguing transport properties [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48].
@@
II Setup of altermagnet-ferromagnet junctions
We consider an AM-FM junction by coupling a two-dimensional AM with a semi-infinite FM lead as depicted in Fig. 1. Region $z<0$ is the FM lead and $z=0$ is the interface. We treat the AM-FM junction as an open quantum system and model it with an effective Hamiltonian of the form,
@@ II Setup of altermagnet-ferromagnet junctions
where $H_{\text{AM}}$ denotes the Hamiltonian of the considered AM system (discussed later), and $\Sigma_{L}$ represents the self-energy arising from the semi-infinite FM lead. In the wide-band limit, the self-energy term becomes independent of both momentum and frequency, and can be analytically calculated as [49, 50, 21, 22],
@@
III Exceptional points in d-wave altermagnets
Let us consider the setup with an FM lead attached to a $d$-wave AM, whose Fermi surface is oriented as shown in Fig. 1(a). The Hamiltonian for a $d$-wave AM is given by [24]
@@ III Exceptional points in d-wave altermagnets
$\gamma^{2}=\lambda^{2}(k_{x}^{2}+k_{y}^{2})+4t_{j1}k_{x}^{2}k_{y}^{2},\quad{\color[rgb]{0,0,0}t_{j1}}\gamma k_{x}k_{y}=0.$ (5)
@@
For the remainder of the paper, for simplicity, we will set $t=1$, $t_{j1}=1$, $\gamma=1$, and $\lambda=1$. The real and imaginary parts of the eigenvalues are plotted in Fig. 2 (a) and (b), respectively. The red dots indicate the merging of the eigenvalues at $(0,\pm 1),(\pm 1,0)$. We further confirm the coalescing of the eigenvectors at the same points by calculating the phase rigidity, $r=\frac{\langle\Psi_{L}|\Psi_{R}\rangle}{\langle\Psi_{R}|\Psi_{R}\rangle}$ [7]. Here $\Psi_{L}$ and $\Psi_{R}$ denote the left and right eigenvectors respectively. Due to bi-orthogonalization, $r\rightarrow 0$ near EPs and approaches unity away from them. In Fig. 2(c) and (d), we show $r$ along $k_{x}$ and $k_{y}$, respectively. We note that $r$ vanishes at $(0,\pm 1)$ and (±$(\pm 1,0)$, which confirms that not only the eigenvalues merge but also the eigenstates coalesce.
As another characterization of the emergence of EPs, we calculate the vorticity, $\nu_{mn}$, given by [51]
@@
Here $E_{m}(\mathbf{k})$ and $E_{n}(\mathbf{k})$ denote the energy eigenvalues in the complex-energy plane and $C$ is a closed loop in the momentum space. In Fig. 2(e), we plot the vorticity of one of the EPs [located at $(0,+1)$] by choosing $C$ around it. We clearly observe the braiding of bands, a direct signature of the underlying EP. This further confirms the appearance of EPs. To examine the order of EPs, we plot the phase rigidity in Fig. 2(f). The blue dots indicate the calculated values of $r$, while the solid green line is the fitted function. The fitting revealed that the exponent of scaling of $r$ is $\sim$0.47, which is close to that expected for a second-order EP [52, 53]. If we take into account the full tight binding model, the earlier results still hold true (for details see Appendix A). Having seen the emergence of EPs in AM-FM junctions, next, we will show how applying a magnetic field allows us to tune and control them.
@@
IV Magnetic field tuning of exceptional points
We next propose and show that the positions of the EPs can be controlled by applying an external magnetic field, $B$, along the interface plane. We find that $B$ can not only change the EP positions asymmetrically, but also merge them beyond a critical value. For illustration, we include the external magnetic field, in our $d$-wave AM-FM junction, along $x$ direction, as shown in Fig. 1. Then, the Hamiltonian in Eq. 3 is augmented by a term of the form $B\sigma_{x}$. The conditions for EPs now read,
$\gamma^{2}=\lambda^{2}k_{x}^{2}+(B+\lambda k_{y})^{2}+4t_{j1}k_{x}^{2}k_{y}^{2},\quad{\color[rgb]{0,0,0}t_{j1}}\gamma k_{x}k_{y}=0.$ (7)
@@ IV Magnetic field tuning of exceptional points
We observe that the position of the EPs can not only be tuned along $k_{x}$ and $k_{y}$ (using $B$), but also can be rotated about the origin by choosing different orientations of the AM Fermi surface with respect to the FM lead. For instance, a $\pi/4$ rotation is achieved with the AM Fermi surface shown in Fig. 1(b). This can be obtained by choosing different AM-specific spin-dependent hopping terms in the Hamiltonian (for details see Appendix B). We, furthermore, discovered that the positions of the emergent EPs can be adjusted to any arbitrary angle. We examine the generalized $d$-wave AM-FM junction [33], characterized by the effective NH Hamiltonian,
$\tilde{H}=$ $H_{\text{AM}}^{d}+\Sigma_{L}(0)$ (8)
$=t(k_{x}^{2}+k_{y}^{2})\sigma_{0}+2t_{j1}k_{x}k_{y}\sigma_{z}+2t_{j2}(k_{x}^{2}-k_{y}^{2})\sigma_{z}$
$+\lambda(k_{y}\sigma_{x}-k_{x}\sigma_{y})+\Sigma_{L}.$
Note that here terms proportional to both $t_{j1}$ and $t_{j2}$ are included in the Hamiltonian. For this case the conditions for obtaining EPs are given by,
$\gamma^{2}=\lambda^{2}(k_{x}^{2}+k_{y}^{2}),\quad\gamma(t_{tj1}k_{x}k_{y}+t_{j2}(k_{x}^{2}-k_{y}^{2}))=0.$ (9)
From Eq. 9, we find that the rotation angle, $\theta=\tan^{-1}(\frac{k_{y}}{k_{x}})$, depends on AM-specific spin-dependent hopping parameters as follows,
$\tan\theta=\frac{-t_{j1}\pm\sqrt{t_{j1}^{2}+4t_{j2}^{2}}}{-2t_{j2}}.$ (10)
By a judicious choice of $t_{j1}$ and $t_{j2}$, we can obtain the EPs along a suitable angle in the $k_{x}-k_{y}$ plane.
@@
V Exceptional points in $g$- and $i$-wave altermagnets
We further extend our analysis to AMs with higher harmonic spin-split Fermi surfaces. Let us first consider a $g$-wave AM-FM junction. The Hamiltonian for a $g$-wave AM reads [24]
@@ V Exceptional points in $g$- and $i$-wave altermagnets
Strikingly distinct from the $d$-wave case, we find that a total of eight EPs emerge at such a junction. The location of EPs are given by $(0,\pm\frac{\gamma}{\lambda})$, (±$(\pm\frac{\gamma}{\lambda},0)$, (±$(\pm\frac{\gamma}{\sqrt{2}\lambda},\pm\frac{\gamma}{\sqrt{2}\lambda})$. These are shown by red dots in Fig. 5(a) and (b). We can extend this analysis to $i$-wave AMs (see Appendix C for details of the Hamiltonian and location of EPs). In this case we find twelve EPs emerging at the AM-FM junction. The real and imaginary parts of the energy eigenvalues are presented in Fig. 5(c) and (d), respectively. We note that on setting $t_{j1}=0$, the $g$-wave case reduces to the $d$-wave case. This is true for $i$-wave scenario as well. In fact, a non-zero hopping term $t_{j1}$ is essential to account for the effect of altermagnet symmetries on the occurrence of EP locations, which results in different number of EPs for different kinds of altermagnets. Therefore, this term is essential to capture the correct physics of the junction. We observe that, in general, there appear twice the EPs as the spin group integer of the AM [23]. In case of bulk altermagnets, augmented by imaginary term proportinal to $\sigma_{z}$, we find that exceptional lines emerge. Their number is twice the number of spin group integer of the altermagnet. We note that the emergent EPs also follow the symmetry of the considered altermagnet. We further note that an applied magnetic field should allow us to tune the EPs in these $g$- and $i$-wave AM-FM junctions, analogous to the $d$-wave case.
VI Summary and outlook
@@
We have proposed and demonstrated the emergence of NH physics, particularly EPs, in AM-FM junctions. We have revealed the presence of EPs in $d$-wave AM-FM junctions and shown their tunability using an external magnetic field. Our findings reveal that EPs can annihilate at a critical value of $B$. Additionally, we have found the tunability of EPs by varying the orientation of the AM Fermi surfaces. We discovered that the number of emergent EPs depends on the spin group integer of the underlying AM. Here we have focused on the EPs in AM-based NH systems. We note that our phase rigidity predictions should be clearly testable, as has been previously measured in several other systems [54, 9]. Furthermore, other features of exceptional points, such as their distinctive energy scaling and concomitant Berry phases [55, 56, 57], would also be present in our proposed platform and could be directly measured in experiments. Furthermore, the non-Hermitian spectra in altermagnet-ferromagnet junctions may be directly amenable to surface spectroscopy measurements, such as angle resolved photoemission spectroscopy. In future, it could be worth exploring other NH aspects of altermagnetic systems, including NH skin effects, NH topology, and interplay with various types of dissipation mechanisms.
Acknowledgments
@@
We thank A. Banerjee, A. Bose, and S. Debadatta for useful discussions. M.A.R. is supported by a graduate fellowship of the Indian Institute of Science. A.N. acknowledges support from the DST MATRICS grant (MTR/2023/000021).
Appendix A:Exceptional points in tight binding model of $d$-wave altermagnets
Here we consider the full tight-binding model for the altermagnet with the self energy arising from the ferromagnetic lead. The full tight binding model with self energy term is given by,
$\tilde{H}=$ $t(\cos^{2}k_{x}+\cos^{2}k_{y})\sigma_{0}+2t_{j1}\sin k_{x}\sin k_{y}\sigma_{z}$ (A1)
$+\lambda(\sin k_{y}\sigma_{x}-\sin k_{x}\sigma_{y})-i\gamma\sigma_{z}.$
The corresponding EP conditions are obtained to be,
$\gamma^{2}=\lambda^{2}(\sin^{2}k_{x}+\sin^{2}k_{y})+4t_{j1}^{2}\sin^{2}k_{x}\sin^{2}k_{y},$ (A2)
$\gamma t_{j1}\sin k_{x}\sin k_{y}=0,$
which give the EP locations at $(0,\pm\sin^{-1}(\gamma/\lambda))$, (±$(\pm\sin^{-1}(\gamma/\lambda),0)$. Therefore, our results are indeed valid considering the full tight-binding model.
Figure 6: EPs in $d$-wave AM-FM junction with $\pi/4$ rotation. (a) Real and (b) imaginary parts of the energy eigenvalues. At the red dots both the eigenvalues merge. Phase rigidity, $r$, along (c) $k_{x}$ and (d) $k_{y}$, respectively. Note that it goes to zero at ($\frac{1}{\sqrt{2}},\pm\frac{1}{\sqrt{2}})$, (±$\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})$ and reaches one away from these coordinates. This indicates the emergence of EPs. We note that the EPs lie along $k_{x}=k_{y}$ and $k_{x}=-k_{y}$, i.e., they are rotated by an angle of $\pi/4$ compared to the $d$-wave case presented in the main text. (e) Vorticity around the EP at $(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})$. The braiding of the bands confirms the emergence of the EPs. (f) Scaling of phase rigidity around the EP. Blue dots are the calculated values and the solid green line represents the fitted function. From the fitted function we obtain an exponent of $\sim$0.46, close to that expected for a second-order EP. We choose $t=1$, $t_{j2}=1$, $\lambda=1$ and $\gamma=1$.
Appendix B: Tuning exceptional points in $d$-wave altermagnets
We can tune the positions of EPs in $k$ space by altering the Fermi surface of the AM as shown in Fig. 1(b) of the main text, which is rotated by an angle of $\pi/4$ with respect to the Fermi surface shown in Fig. 1(a) of the main text. This rotates the EP positions by an angle of $\pi/4$. This can be achieved by choosing a different $d$-wave AM-FM junction effective NH Hamiltonian, which reads as follows,
$\tilde{H}=$ $H_{\text{AM}}^{d}+\Sigma_{L}(0)$ (A3)
$=t(k_{x}^{2}+k_{y}^{2})\sigma_{0}+2t_{j2}(k_{x}^{2}-k_{y}^{2})\sigma_{z}$
$+\lambda(k_{y}\sigma_{x}-k_{x}\sigma_{y})+\Sigma_{L}.$
Here $t_{j2}$ is AM-specific spin-dependent hopping parameter distinct from $t_{j1}$. The other terms in the Hamiltonian are as already defined in the main text. Following similar calculations as in the main text, the conditions for the emergence of EPs turn out to be,
$\gamma^{2}=\lambda^{2}(k_{x}^{2}+k_{y}^{2})+4t_{j2}^{2}(k_{x}^{2}-k_{y}^{2})^{2},\quad\gamma t_{j2}(k_{x}^{2}-k_{y}^{2})=0.$ (A4)
From Eq. A4, we find that the four EP positions are ($\frac{\gamma}{\sqrt{2}\lambda}$,±$\pm\frac{\gamma}{\sqrt{2}\lambda}$) and (±$\pm\frac{\gamma}{\sqrt{2}\lambda}$,$\frac{\gamma}{\sqrt{2}\lambda}$). So, the EP positions not only come closer to the origin, but are also rotated by an angle of $\pi/4$, as compared to the $d$-wave AM whose Fermi surface orientation is as shown in panel Fig. 1(a) of the main text. We set $t=1$, $t_{j2}=1$, $\lambda=1$ and $\gamma=1$, which leads to four EPs at ($\frac{1}{\sqrt{2}},\pm\frac{1}{\sqrt{2}})$, (±$\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})$ in the $k_{x}-k_{y}$ plane.
We further confirm the emergence of EPs by plotting the real and imaginary parts of the energy eigenvalues in Fig. 6(a) and (b), respectively. We observe that at the red dots, the eigenvalues merge. We further investigate the phase rigidity $r$ in Fig. 6(c) and (d) along $k_{x}$ and $k_{y}$, respectively. We note that $r$ takes the value of zero at the positions of the red dots (EPs), as expected. In Fig. 6(e), we plot the vorticity around the EP and the interchange of bands confirms the emergence of the EP. We further investigate the scaling of the phase rigidity near the EP. In Fig. 6(f) the blue dots show the calculated $r$ values and the solid green line is the fitted function. The fitted function yields the exponent to be $\sim$0.46, which is close to the exponent expected for a second-order EP [52, 53].
Appendix C: Exceptional points in $i$-wave altermagnets
We consider an $i$-wave AM [24]. The corresponding AM-FM junction is described by the following NH Hamiltonian,
$H_{\text{AM}}^{i}=$ $t(k_{x}^{2}+k_{y}^{2})\sigma_{0}+2t_{j}k_{x}k_{y}(3k_{x}^{2}-k_{y}^{2})(3k_{y}^{2}-k_{x}^{2})\sigma_{z}$ (A5)
$+\lambda(k_{y}\sigma_{x}-k_{x}\sigma_{y})+\Sigma_{L}.$
Following similar calculations as the $d$- and $g$-wave cases, we find the following conditions for the emergence of EPs,
$\gamma^{2}=\lambda^{2}(k_{x}^{2}+k_{y}^{2})+4t_{j}^{2}k_{x}^{2}k_{y}^{2}(3k_{x}^{2}-k_{y}^{2})^{2}(3k_{y}^{2}-k_{x}^{2})^{2},$ (A6)
$\gamma t_{j}k_{x}k_{y}(3k_{x}^{2}-k_{y}^{2})(3k_{y}^{2}-3k_{x}^{2})=0.$
From Eq. A6, we obtain the location of the EPs as $(0,\pm\frac{\gamma}{\lambda})$, (±$(\pm\frac{\gamma}{\lambda},0)$, (±$(\pm\frac{\gamma}{2\lambda},\pm\frac{\sqrt{3}\gamma}{2\lambda})$, (±$(\pm\frac{\sqrt{3}\gamma}{2\lambda},\pm\frac{\gamma}{2\lambda})$. We note that a total of twelve EPs emerge for the $i$ wave AM-FM junction. These are indicated by red points in Fig. 5(c) and (d) of the main text.